OpenStudy (anonymous):

A hailstone traveling with a velocity of 43 meters/second comes to a virtual stop 0.28 seconds after hitting water. What is the magnitude of its acceleration in the water?

OpenStudy (badhi):

use $$v=u+at$$

OpenStudy (anonymous):

Acceleration is the change in velocity divided by time. So \[\frac{ Vf - Vi }{ T }\] Vf = 0 and Vi = 43 and T = .28s So \[a = \frac{ 0 - 43 }{ .28 }\]